The differential equation which has y 4 = Cx3 − 3 as its general solution is:


1
Expert's answer
2022-01-28T09:20:59-0500

"y^ 4 = Cx^3 \u2212 3" ...(1)

Differentiating both sides w.r.t "x" , we get

"4y^3y'=3Cx^2\Rightarrow C=\frac {4y^3y'}{3x^2}"

Put the value of C in equation (1), we get

"y^4=\frac {4y^3y'}{3x^2}\times x^3-3=\frac {4xy^3y'}{3}-3=\frac {4xy^3y'-9}{3}\\
3y^4-4xy^3y'+9=0"

is the differential equation.